# Large Scale White Noise and Cosmology

### v1

@@ I.4 Outline
    provides analytical expressions for the LSWN kernel functions $\mathcal{M}_{\pm}[\varphi]$ in self-similar systems with quadratic non-linearities. These functions, expressed in terms of regularized hypergeometric functions, determine the amplitude of LSWN generated by mode coupling and exhibit characteristic power-law behavior at small and large scales with a transition ("knee") at horizon crossing.
@@ II.1.2 The Conservative Exception
Other than these exceptions, we expect all sshPDEs to exhibit LSWN. Now, one can always reparameterize a PDE via a change of variables (with corresponding changes to the initial conditions) in order to simplify the system. [^7] We refer to the PDE with the statistical distribution of initial conditions as the system. This includes reparameterizations that transform non-linear PDEs, which feature LSWN, into either linear or conservative PDEs, which exhibit no LSWN. [^8] For example, one can transform any system exhibiting LSWN into another system exhibiting no LSWN using the transformation $q[t,\vec{x}]\rightarrow\underline{q}[t,\vec{x}]=\nabla^{2}q[t,\vec{x}]$. In this sense, one may worry that the amount or even existence of LSWN is tied to a particular parametrization of the PDE, and therefore might seem mathematically ill-defined.
@@ III Large Scale Noise in Self-Similar Systems
where we are again using the sound horizon $h$ as the temporal variable. The first line gives the $m=1$ (linear) order differential operator, and the second line gives the $m=2$ order operator, etc. Assume that $\zeta_{3},\zeta_{4},\zeta_{5},\zeta_{6}\,\zeta_{7}\lesssim 1$ and at least one of these is non-zero. In this case, the leading order non-linearity is quadratic, i.e. $n=2$.
@@ III.1 Late-Time Relic Systems
$P_{(2)}[h,k]\approx P_{(2)}^{\mathrm{lin}}[h,k]+P_{(2)}^{\mathrm{nl}}[h,k]\,,\qquad P_{(2)}^{\mathrm{lin}}[h,k]=R_{(2)}[k]\,{\cal Q}_{+}[k\,h]^{2}\,,\qquad P_{(2)}^{\mathrm{nl}}[h,k]=R_{(2)}^{\mathrm{nl}}[k]\,{\cal Q}_{+}[k\,h]^{2}\,$, (141)
where $P_{(2)}^{\mathrm{lin}}$ is the linear theory prediction given the true initial conditions and $P_{(2)}^{\mathrm{nl}}$ is the corrections due to non-linearities. At the longest wavelengths, one has
@@
where $V_{S^{d}}$ is the volume of the unit $d$-sphere. If $\mathfrak{W}_{+}\neq 0$, this is positive and finite so long as the $k_{1}$ integral is finite. This requires $R_{(2)}[k]$ be cutoff at large $k$. If all these conditions are satisfied, then non-linearities add white noise on scales larger than the sound horizon. If the linear theory power spectrum is sub-Poissonian, i.e. if $P_{(2)}^{\mathrm{lin}}[0^{+}]=0$, then the white noise from non-linear corrections will dominate on the largest scales.
@@ III.1.1 Conservative Relics
so the spectral index is 2 rather than 0. Thus, while there is no LSWN, if the spectral index of $R_{(2)}[k]$ is greater than 2, this large scale noise will still dominate at the largest scales.
@@ III.1.2 Cosmic Confusion
The LOBA results of the above subsections give an additive correction, $R_{(2)}^{\mathrm{nl}}[k]$, to the initial power, $R_{(2)}[k]$ on larger scales. If $R_{(2)}[k]$ falls off faster at small $k$ than $R_{(2)}^{\mathrm{nl}}[k]$, then it is the non-linear correction which will dominate. If one only measures $P_{(2)}[k]$ on scales where the non-linear corrections dominate, one might mistakenly conclude that the power observed on these scales represents linear growth of initial conditions with a very different spectrum from the actual initial conditions. One can attribute the power observed on these scales to either initial conditions or non-linearities on scales much smaller than those observed. We refer to this ambiguity as cosmic confusion.
The large scale power spectra we have obtained from LOBA are power laws with spectral index $d\ln[P_{(2)}^{\mathrm{nl}}]/d\ln[k]\approx 0$ (LSWN) unless the non-linearities are conservative, in which case $d\ln[P_{(2)}^{\mathrm{nl}}]/d\ln[k]\approx 2$. We do not expect deviations from these power laws with non-self-similar non-linearities, but proving this will require further investigation. Thus, not observing a power law might be a hint that what one is observing is not a produce of non-linearities.
@@ III.2 Choice of Dependent Variable
for $\underline{q}$. Since the two parameters characterizing the linear EoM, $\zeta+1$ and $\zeta_{2}$ the growing and decaying modes, $\cal Q$$_{\pm}$ are unchanged. Three of the five parameters characterizing the non-linear part of the EoM are changed if $\alpha\neq 0$. The LOBA normalization of the non-conservative late time relics is also changed
@@
From this we see that there is an $\alpha$ and therefore choices of $\underline{q}$ where $\underline{\mathfrak{W}}_{+}=0$. While generically $\mathfrak{W}_{+}\sim{\cal O}[1]$ it can be very small or zero even zero. For a given physical system the measured amplitude of the LSWN which is $\propto{\underline{\mathfrak{W}}_{+}}^{2}$ may or may not be very small depending on what one is measuring. This is another manifestation of the tyranny of observables (§II.1.2).
@@ III.3 Perturbative Initial Conditions
The dynamical systems we are modeling are reversible. Thus any final state can be obtained by appropriate choice of initial conditions. All claims we have made about the existence or non-existence of LSWN are dependent on our assumptions about these initial conditions. The assumption we have made in much of this paper is that the initial conditions are pure growing mode initial conditions, which we review below.
@@
It is important to emphasize here that the initial conditions are a choice not determined by the EoMs. When describing a physical system one must use physics to decide the appropriate initial conditions. The three types of initial conditions for perturbation theory described in the previous paragraph are not exhaustive and the appropriate initial condition might be something very different. There are many choices among the $p\times n$ parameters for each mode. Applications of this formalism of most interest here are for systems which are statistical ensembles of initial conditions. One should distinguish the properties of initial conditions of a particular instance and properties of all instances of the ensemble which adds another level of complication.
Choices of dependent variables and of initial conditions are major determinants of the manifestations of non-linearities. A more comprehensive presentation of this topic will be given in Ref. [5].
@@ IV.1 Lagrangian Coordinates
Parameterizing the inhomogeneity using $\delta$ is the conventional choice, using $\delta\lambda$ has the nice feature that the EoM is non-singular for all finite values, and using $\mathfrak{d}$ yields the simplest EoM. To linear order in amplitude, these are all the same: $\delta+{\cal O}[\delta^{2}]=\delta\lambda+{\cal O}[\delta\lambda^{2}]=\mathfrak{d}+{\cal O}[\mathfrak{d}^{2}]$. To higher order, however, they differ. One can in principle use any of these parameters as dependent variables because they are perturbation variables, meaning zero ($\rho=\bar{\rho}$) is a solution to the equations of motion. They thus obey homogeneous PDEs. Since these different dependent variables are not linearly related to each other, they are deformations of each other (see §II.1.4) and will have different levels of LSWN.
@@ IV.2 Matter Era
where $\mathsf{g}[\chi]$ is the local amplitude of the growing mode and $\mathsf{d}[\chi]$ is the local amplitude of the decaying mode. This solution is only valid so long as $\mathfrak{d}<1$, at which point caustics form. Thus, the solution extends to $t=0$ only if $\mathsf{d}[\chi]\leq 0\ \forall\chi$. For this reason we will henceforth only consider pure growing mode solutions: $\mathsf{d}[\chi]=0$. Even with purely growing modes, the solution is only valid in the finite time interval $0\leq t<t_{\mathrm{fid}}/\max_{\chi}[\mathsf{d}[\chi]]^{3/2}$.
@@
for small $\mathsf{g}$ but any $t$. This $2^{\rm nd}$ order expansion is exactly the LOBA Eq. (126) given by the parameters from Table 1. The non-linearities remain active (see §II.2.1) since the non-linear correction grows $\propto t^{4/3}$ and not $\propto t^{2/3}$, as predicted by linear theory. The $2^{\rm nd}$ order corrections remain small and the Taylor series remains accurate when and where $t\ll t_{\mathrm{fid}}\,|\mathsf{g}[\chi]|^{-3/2}$, i.e. while in the linear regime, $|\delta|,\,|\delta\lambda|,\,|\mathfrak{d}|\ll 1$.
@@ IV.3 Radiation Era
The factors of order unity have to do with the shape of the cutoff and the precise definition of $\sigma$. These could vary significantly but the general scaling with $n_{\mathrm{s}}$ should hold so long at the cutoff is sufficiently sharp. When the inhomogeneities are very small ($A\ll 1)$ the cutoff scale is much smaller than this transition scale and it might be impossible determine $\sigma$ directly. Eq. 169 allows one to use the large scale inhomogeneities as a "microscope" to infer things about very small scale inhomogeneities.
@@
According to linear theory the rms overdensity increases while the inhomogeneities are outside the horizon and after crossing the horizon when $h\sim h_{*}\sim{k_{*}}^{-1}$ the inhomogeneities are Hubble damped and the rms overdensity decreases. At all times $\langle\delta^{2}\rangle\ll 1$ in linear theory so one might naively think that linear theory is always accurate This is not true! The actual evolution does follow the linear timeline, peaking at $\delta_{\mathrm{peak}}$ until long past horizon crossing of the initial perturbations. However during horizon crossing a small amplitude long wavelength tail is generated by non-linearities. These wavelengths remain outside the sound horizon, are not damped, but continue to grow. Eventually, at a time well after horizon crossing, the contribution of the growing long wavelength inhomogeneities to $\langle\delta^{2}\rangle$ will exceed the linear theory prediction. This is a time when one might say that linear theory "breaks down". Once the long wavelengths start to dominate $\langle\delta^{2}\rangle$ will continue to grow and eventually, very long after breakdown the inhomogeneities will "go non-linear" i.e. $\langle\delta^{2}\rangle\sim 1$. At this point the flow will be supersonic and shocks will form. As the fluid begins to go non-linear the LOBA will become increasingly inaccurate. In this paper we do not address the question of how the fluid evolves after it goes non-linear.
@@
It is instructive to consider quantitative aspects this timeline. For parameters used for Fig. 2 the rms overdensity peaks at $\delta_{\mathrm{peak}}\sim 10^{-5}$ when the comoving horizon, $h=h_{*}$, equals the comoving wavelength of the initial inhomogeneities. Much later linear theory breaks down when the comoving horizon is $\sim 10^{2}\,h_{*}$. Much much later the fluid goes non-linear. The comoving wavelengths which first go non-linear are $\sim 10^{12}$ times larger than the comoving wavelengths of the initial inhomogeneities! The range of temporal scales is even more striking. Using $h=3\,c_{\mathrm{s}}\,{t_{\mathrm{fid}}}^{2/3}\,t^{1/3}$ if $t_{*}$ is the time of horizon crossing then the time when linear theory breaks down is $10^{6}\,t_{*}$ and the time when the fluid goes non-linear is $10^{36}\,t_{*}$. The range of spatial and temporal scales would be even larger for smaller $\delta_{\mathrm{peak}}$. No matter how small the initial inhomogeneity non-linearity will eventually dominate the dynamics however one must observe a very larger range of timescales and length scales see these effects. Non-linear effects are not easily discerned when on considers a small range of scales over a limited time. Even though cosmic inhomogeneities are initially very small the universe is a very big place which has existed for a very long time so cosmology is system where non-linear effects will likely manifest.
@@
Figure 2: For a range of times parameterized by the sound horizon, $h$ as defined in Eq. 161) we plot the estimated power spectrum from Eq. IV.3 for $\delta$ in blue and the linear theory prediction in green. Initial conditions are of the form of Eq. 166 with $n_{\mathrm{s}}=6$ and normalized such that $\delta_{\mathrm{rms}}^{\mathrm{linear}}$ peaks at $10^{-5}$. All quantities are normalized to the cutoff scale, $\sigma$. Plotted in green is the linear theory prediction. Evident are the acoustic oscillations on small scales and as well as large scale white noise (LSWN). Linear theory evolution dominates initially while non-linearities generate additional large scale power which eventually also grows like linear theory, i.e. growth outside the sound horizon and damping inside.

### v2

@@ I.4 Outline
    provides analytical expressions for the LSWN kernel functions $\mathcal{M}_{\pm}[\varphi]$ in self-similar systems with quadratic non-linearities. These functions, expressed in terms of regularized hypergeometric functions, determine the amplitude of LSWN generated by mode coupling and exhibit characteristic power-law behavior at small and large scales with a transition (“knee”) at horizon crossing.
@@ II.1.2 The Conservative Exception
To be clear the quantity that must be conserved is $q$, it does not matter if there is some other quantity which is conserved. Furthermore since we are concerned with the power spectrum of the conventional Fourier transform in Euclidean $\vec{x}$ the conservations law is that the Euclidean volume weighted $q$ is conserved. It does not matter if $q$ is conserved with some other weighting.
Other than these exceptions, we expect all sshPDEs to exhibit LSWN. These are exceptions only if the initial conditions are sub-Poissonian which, as emphasized in §III.3, will often not be the case. Now, one can always reparameterize a PDE via a change of variables (with corresponding changes to the initial conditions) in order to simplify the system. [^7] We refer to the PDE with the statistical distribution of initial conditions as the system. This includes reparameterizations that transform non-linear PDEs, which feature LSWN, into either linear or conservative PDEs, which exhibit no LSWN. [^8] For example, one can transform any system exhibiting LSWN into another system exhibiting no LSWN using the transformation $q[t,\vec{x}]\rightarrow\underline{q}[t,\vec{x}]=\nabla^{2}q[t,\vec{x}]$. In this sense, one may worry that the amount or even existence of LSWN is tied to a particular parametrization of the PDE, and therefore might seem mathematically ill-defined.
@@ II.2.2 Growing Mode Initial Conditions
In a perturbative expansion initial conditions are also perturbative (see §III.2) and this growing mode prescription is to all orders, which we call pure growing mode initial conditions. For $\mathcal{O}[n]$ initial conditions it only matters that this is true to $\mathcal{O}[n]$. Any $q$ that has this type of initial condition we call a guiding variable for reasons described in §III.3. There it is shown that if $q$ as guiding variable then any for deformation $\underline{q}=F[q]$ non-linear at $\mathcal{O}[n]$ that if $q$ is a guiding variable then $\underline{q}$ cannot be. Guiding variables are special and must be carefully chosen for your application.
@@ II.6 Review
  • growing mode initial conditions and guiding variables
@@ III Large Scale Noise in Self-Similar Systems
where we are again using the sound horizon $h$ as the temporal variable. The first line gives the $m=1$ (linear) order differential operator, and the second line gives the $m=2$ order operator, etc. Assume that $|\zeta_{3}|,|\zeta_{4}|,|\zeta_{5}|,|\zeta_{6}|\,|\zeta_{7}|\lesssim 1$ and at least one of these is non-zero. In this case, the leading order non-linearity is quadratic, i.e. $n=2$.
@@ III.1 Late-Time Relic Systems
$P_{(2)}[h,k]\approx P_{(2)}^{\mathrm{lin}}[h,k]+P_{(2)}^{\mathrm{nl}}[h,k]\,$, (141)
where
$P_{(2)}^{\mathrm{lin}}[h,k]=R_{(2)}[k]\,{\cal Q}_{+}[k\,h]^{2}\,,\qquad P_{(2)}^{\mathrm{nl}}[h,k]=R_{(2)}^{\mathrm{nl}}[k]\,{\cal Q}_{+}[k\,h]^{2}\,.$ (142)
Here $P_{(2)}^{\mathrm{lin}}$ is the linear theory prediction given the true initial conditions and $P_{(2)}^{\mathrm{nl}}$ is the corrections due to non-linearities. At the longest wavelengths, one has
@@
where $V_{S^{d}}$ is the volume of the unit $d$-sphere. If $\mathfrak{W}_{+}\neq 0$, this is positive and finite so long as the $k_{1}$ integral is finite. This requires $R_{(2)}[k]$ be cutoff at large $k$. If all these conditions are satisfied, then non-linearities add white noise on scales larger than the sound horizon. If the linear theory power spectrum is sub-Poissonian, i.e. if $P_{(2)}^{\mathrm{lin}}[0^{+}]=0$, then this white noise from non-linear corrections will dominate on the largest scales.
@@ III.1.1 Conservative Relics
so the spectral index is 2 rather than 0. Thus, while there is no LSWN, if the spectral index of $R_{(2)}[k]$ is greater than 2, this large scale (blue) noise will still dominate at the largest scales.
@@ III.1.2 Cosmic Confusion
The LOBA results of the above subsections give an additive correction, $R_{(2)}^{\mathrm{nl}}[k]$, to the initial power, $R_{(2)}[k]$ on larger scales. If $R_{(2)}[k]$ falls off faster at small $k$ than $R_{(2)}^{\mathrm{nl}}[k]$, then it is the non-linear correction which will dominate. If one only measures $P_{(2)}[k]$ on scales where the non-linear corrections dominate, one might mistakenly conclude that the power observed on these scales represents linear growth of initial conditions with a very different spectrum from the actual initial conditions. Thus one can attribute the power observed on these scales to either initial conditions or non-linearities on scales much smaller than those observed. We refer to this ambiguity as cosmic confusion.
The large scale power spectra we have obtained from LOBA are power laws with spectral index $d\ln[P_{(2)}^{\mathrm{nl}}]/d\ln[k]\approx 0$ (LSWN) unless the non-linearities are conservative, in which case $d\ln[P_{(2)}^{\mathrm{nl}}]/d\ln[k]\approx 2$. We do not expect deviations from these power laws with non-self-similar non-linearities, but proving this will require further investigation. Thus, not observing a power law might be a hint that what one is observing is not a product of non-linearities.
@@ III.2 Choice of Dependent Variable
for $\underline{q}$. Since the two parameters characterizing the linear EoM, $\zeta_{1}$ and $\zeta_{2}$ the growing and decaying modes, $\cal Q$$_{\pm}$ are unchanged. Three of the five parameters characterizing the non-linear part of the EoM are changed if $\alpha\neq 0$. The LOBA normalization of the non-conservative late time relics is also changed
@@
From this we see that there is usually an $\alpha$ and therefore choices of $\underline{q}$ where $\underline{\mathfrak{W}}_{+}=0$. We call such a suppressed LSWN variable. While generically one would expect that $\mathfrak{W}_{+}\sim{\cal O}[1]$ we now seed that it can be very small or zero even zero. For a given physical system the measured amplitude of the LSWN, $\propto{\underline{\mathfrak{W}}_{+}}^{2}$, may or may not be very small depending on what one is measuring. This is another manifestation of the tyranny of observables (§II.1.2).
@@ III.3 Perturbative Initial Conditions
The dynamical systems we are modeling are reversible. Thus any final state can be obtained by appropriate choice of initial conditions. All claims we have made about the existence or non-existence of LSWN are dependent on our assumptions about these initial conditions. The assumption we have made in much of this paper is that the initial conditions are pure growing mode initial conditions, which we now review.
@@
It is important to emphasize here that the initial conditions are a choice not determined by the EoMs. When describing a physical system one must use physics to decide the appropriate initial conditions. The three types of initial conditions for perturbation theory described in the previous paragraph are not exhaustive and the appropriate initial condition might be something very different. There are many choices among the $p\times n$ parameters for each mode. Applications of this formalism of most interest here are for systems which are statistical ensembles of initial conditions. One should distinguish the properties of initial conditions of a particular instance and properties of all instances of the ensemble. This adds another level of complexity.
III.3.1 Guiding Variables at the Initial Singularity
The above prescription becomes is obscure when the initial conditions are at an initial singularity. This is the case of most interest to us! One might erroneously think that when the growing mode becomes arbitrarily small at the initial singularity that the linear theory becomes more and more accurate and therefore one need only set the initial conditions at linear order. This might lead one to believe that all choices of perturbation variable that are equal at linear order, e.g. $\underline{q}=q-\alpha\,q^{2}+\cdots$, have the same initial conditions. Furthermore one might conclude that all $\underline{q}$ can have sub-Poissonian initial conditions. All this is wrong for the simple reason that increasingly small difference in solutions as one approaches the singularity are amplified by subsequent growth and not negligible.
To see this we may use the fact that the gradient terms in Eq. (131) become irrelevant as one approach the initial singularity at $h=0$. This is a property of self-similar systems. In this limit, Eq. (131) reduces to the hyperlocal equations considered in §II.5. As noted there, by change of the temporal variable $t\rightarrow\mathfrak{t}$ and dependent variable $q\rightarrow\mathfrak{q}$, one can convert to an equivalent system in canonical form, where the linear order operator is $_{(1)}\hat{\mathcal{D}}\,\mathfrak{q}=\ddot{\mathfrak{q}}[\mathfrak{t},\vec{x}]$. The canonical form will differ for deformations of the dependent variable, $\mathfrak{q}\rightarrow\underline{\mathfrak{q}}=\mathfrak{q}-\alpha\,\mathfrak{q}^{2}+\cdots$, but the linear operator remains the same. Thus, both $\mathfrak{q}$ and $\underline{\mathfrak{q}}$ have growing and decaying modes $\mathfrak{t}$ and $1$. If $\mathfrak{q}$ has pure growing mode initial conditions, then near the singularity $\mathfrak{q}[\mathfrak{t},\vec{x}]=\mathfrak{q}_{0}[\vec{x}]\,\mathfrak{t}+\mathcal{O}[\mathfrak{t}^{3}]$ while $\underline{\mathfrak{q}}$ exhibits very different behavior, $\underline{\mathfrak{q}}[\mathfrak{t},\vec{x}]=\mathfrak{q}_{0}[\vec{x}]\,\mathfrak{t}-\alpha\,\mathfrak{q}_{0}[\vec{x}]^{2}\,\mathfrak{t}^{2}+\mathcal{O}[\mathfrak{t}^{3}]$. The $\mathfrak{q}_{0}^{2}$ terms is the 2nd order part of the initial condition for $\underline{\mathfrak{q}}$ that must not be neglected. If $\alpha\neq 0$ there are at least three properties in which $\mathfrak{q}$ and $\underline{\mathfrak{q}}$ must differ:
  1. 1.
    $\mathfrak{q}$ and $\underline{\mathfrak{q}}$ cannot both be guiding variables,
  2. 2.
    $\mathfrak{q}$ and $\underline{\mathfrak{q}}$ cannot both be conservative,
  3. 3.
    $\mathfrak{q}$ and $\underline{\mathfrak{q}}$ cannot both have sub-Poissonian initial conditions
because
  1. 1.
    $\underline{\mathfrak{q}}$ is not a guiding variable due to the quadratic term just mentioned,
  2. 2.
    a conservative quantity must have $\zeta_{3}=0$, and $\alpha\neq 0$ means that $\zeta_{3}\neq\underline{\zeta}_{3}$,
  3. 3.
    since the square of a sub-Poissonian function is a Poissonian function, the $\mathfrak{q}_{0}[\vec{x}]^{2}\mathfrak{t}^{2}$ difference means that $\ddot{\tilde{\mathfrak{q}}}[0,\vec{0}^{+}]\neq\ddot{\tilde{\underline{\mathfrak{q}}}}[0,\vec{0}^{+}]$.
This has a variety of interesting implications, among them:
  • one must specify among quantities which are identical to 1st order which (if any) have pure growing mode initial conditions to leading non-linear order ($n$), and
  • a quantity that is conservative or has a linear EoM will develop LSWN unless it has both sub-Poissonian initial conditions and pure growing mode initial conditions.
Thus, protection from LSWN hinges on initial conditions too!
Many of our results up to this point in §II-III use the assumption that $q$ is a guiding variable, meaning it has growing mode initial conditions at order 1 and zero initial conditions at order 2 to $n$. For example if $q$ is a guiding variable in the relic systems of §III.1 and if $q$ has sub-Poissonian initial conditions then if $q$ has a non-linear non-conservative EoM and will generate LSWN. One can then compute the power spectrum of the deformation $\underline{q}=F[q]$ using the function $F[q]=q-\alpha q^{2}+\cdots$ to determine the power spectrum for $\underline{q}$. One will find that $\underline{q}$ will also have LSWN whether or not it has a linear or conservative EoM i.e. $\underline{q}$ "inherits" LSWN from $q$. Only the $\mathfrak{W}_{+}$ of the guiding variable really matters. A guiding variable plays a special role in determining the large scale evolution of all of its deformations which is why we use "guiding".
For physical systems, one must know the initial conditions in detail in order to predict how it manifests non-linearities. One must know what are the guiding variables and which have sub-Poissonian initial conditions.
@@ IV.1 Lagrangian Coordinates
Parameterizing the inhomogeneity using $\delta$ is the conventional choice, using $\delta\lambda$ has the nice feature that the EoM is non-singular for all finite values, and using $\mathfrak{d}$ yields the simplest EoM. To linear order in amplitude, these are all the same: $\delta+{\cal O}[\delta^{2}]=\delta\lambda+{\cal O}[\delta\lambda^{2}]=\mathfrak{d}+{\cal O}[\mathfrak{d}^{2}]$. To higher order, however, they differ. One can in principle use any of these parameters as dependent variables because they are perturbation variables, meaning zero ($\rho=\bar{\rho}$) is a solution to the equations of motion. They thus obey homogeneous PDEs. Since these different dependent variables are not linearly related to each other, they are deformations of each other (see §II.1.4) and may have different levels of LSWN. As explained in §III.3 we can choose one of these to be a guiding variables i.e. to have pure growing mode initial conditions, and this means that others will not. The subsequent evolution of the system can be quite different depending on the choice of guiding variable.
@@ IV.2 Matter Era
where $\mathsf{g}[\chi]$ is the local amplitude of the growing mode and $\mathsf{d}[\chi]$ is the local amplitude of the decaying mode. This solution is only valid so long as $\mathfrak{d}<1$, otherwise caustics form. Thus, the solution extends to $t=0$ only if $\mathsf{d}[\chi]\leq 0\ \forall\chi$. To avoid caustics at early times requires $\mathsf{d}[\chi]=0$ in which case $\mathsf{d}$ is a guiding variable and therefore $\delta$ and $\delta\lambda$ cannot be (see §III.3. Furthermore if $\mathfrak{d}$ has sub-Poissonian initial conditions then neither $\delta$ or $\delta\lambda$ can. Even with $\mathsf{d}=0$ one may not evade singular behavior since caustics will form when $t=t_{\mathrm{fid}}/\max_{\chi}[\mathsf{d}[\chi]]^{3/2}$.
@@
for small $\mathsf{g}$ but any $t$. This $2^{\rm nd}$ order expansion is exactly the LOBA of Eq. (126) given by the parameters from Table 1. The non-linearities remain active (see §II.2.1) since the non-linear correction grows $\propto t^{4/3}$ and not $\propto t^{2/3}$, as predicted by linear theory. The $2^{\rm nd}$ order corrections remain small and the Taylor series remains accurate when and where $t\ll t_{\mathrm{fid}}\,|\mathsf{g}[\chi]|^{-3/2}$, i.e. while in the linear regime, $|\delta|,\,|\delta\lambda|,\,|\mathfrak{d}|\ll 1$.
@@ IV.3 Radiation Era
The factors of order unity have to do with the shape of the cutoff which precisely defines $\sigma$. These could vary significantly but the general scaling with $n_{\mathrm{s}}$ should hold so long at the cutoff is sufficiently sharp. When the inhomogeneities are very small ($A\ll 1)$ the cutoff scale is much smaller than this transition scale and it might be impossible determine $\sigma$ directly. Eq. 170 allows one to use the large scale inhomogeneities as a "microscope" to infer things about very small scale inhomogeneities.
@@
According to linear theory the rms overdensity increases while the inhomogeneities are outside the horizon and after crossing the horizon when $h\sim h_{*}\sim{k_{*}}^{-1}$ the inhomogeneities are Hubble damped and $\langle\delta^{2}\rangle$ decreases. At all times $\langle\delta^{2}\rangle\ll 1$ in linear theory so one might naively think that linear theory is always accurate This is not true! The actual evolution does follow the linear timeline, peaking at $\delta_{\mathrm{peak}}$ until long past horizon crossing of the initial perturbations. However during this horizon crossing a small amplitude long wavelength tail is generated by non-linearities. These wavelengths remain outside the sound horizon, are not damped, but continue to grow. Eventually, at a time well after horizon crossing, the contribution of the growing long wavelength inhomogeneities to $\langle\delta^{2}\rangle$ will exceed the linear theory prediction. This is a time when one might say that linear theory “breaks down”. Once the long wavelengths start to dominate $\langle\delta^{2}\rangle$ will continue to grow and eventually, very long after breakdown the inhomogeneities will “go non-linear”, i.e. $\langle\delta^{2}\rangle\sim 1$. At this point the flow will be supersonic and shocks will form. As the fluid begins to go non-linear the LOBA will become increasingly inaccurate. In this paper we do not address the question of how the fluid evolves after it goes non-linear.
@@
It is instructive to consider quantitative aspects this timeline. For parameters used for Fig. 2 the rms overdensity peaks at $\delta_{\mathrm{peak}}\sim 10^{-5}$ when the comoving horizon, $h=h_{*}$, equals the comoving wavelength of the initial inhomogeneities. Much later linear theory breaks down when the comoving horizon is $\sim 10^{2}\,h_{*}$. Much much later the fluid goes non-linear. The comoving wavelengths which first go non-linear are $\sim 10^{6}$ times larger than the comoving wavelengths of the initial inhomogeneities! The range of temporal scales is even more striking. Using $h=3\,c_{\mathrm{s}}\,{t_{\mathrm{fid}}}^{2/3}\,t^{1/3}$ if $t_{*}$ is the time of horizon crossing then the time when linear theory breaks down is $10^{6}\,t_{*}$ and the time when the fluid goes non-linear is $10^{18}\,t_{*}$. The range of spatial and temporal scales would be even larger for shallower initial slope, i.e. $n_{\mathrm{s}}<5$. A steep slope was chosen to avoid a much larger range of scales. No matter how small the initial inhomogeneity non-linearity will eventually dominate the dynamics however one must observe a very large range of timescales and length scales see these effects. Non-linear effects are not easily discerned when one considers a small range of scales over a limited time. Even though cosmic inhomogeneities are initially very small the universe is a very big place which has existed for a very long time so cosmology is a system where non-linear effects have a chance to manifest.
@@
Figure 2: For a range of times parameterized by the sound horizon, $h$ as defined in Eq. (162) we plot the estimated power spectrum from Eq. (IV.3) for $\delta$ in blue and the linear theory prediction in green. Initial conditions are of the form of Eq. (167) with $n_{\mathrm{s}}=5$ and normalized such that $\delta_{\mathrm{rms}}^{\mathrm{linear}}$ peaks at $10^{-5}$. All quantities are normalized to the cutoff scale, $\sigma$. Plotted in green is the linear theory prediction. Evident are the acoustic oscillations on small scales and as well as large scale white noise (LSWN). Linear theory evolution dominates initially while non-linearities generate additional large scale power which eventually also grows like linear theory, i.e. growth outside the sound horizon and damping inside.